determine if the lines are parallel, perpendicular or neither. y=2x + 6 and y=-1/2x+10
Answer by mananth(16946)

y=2x + 6
m1=2
y=-1/2x+10
m2=-1/2
m1*m2= 2*-1/2 =-1
so the lines are perpendicular